Question 20 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes. With specialization, Spain produces only grapes-75 tons-and France produces 75 ships. The terms of trade are 1 ship for 1 ton of grapes, and 30 ships are traded for 30 tons of grapes. Are the countries better off or worse off with trade? With specialization and trade, how many additional ships (if any) can Spain consume? Provide your answer as a number rounded to two decimal places. Use - to indicate negative amounts. Do not include any symbols, such as s. , %, or , in your answer Spain France Shi Grapes (in tons 75 Shi 10 20 30 40 50 Grapes (in tons) 50 40 30 20 10 45 30 15 30 45 0 75
Answer #1

With trade,

Spain exports 30 grapes to France and imports 30 ships from France.

France exports 30 ships to Spain and imports 30 grain from Spain.

After importing 30 ships, Spain consumes (0 + 30) = 30 ships, which is the same as it was producing & consuming before trade.

Increase in consumption of ships = 30 - 30 = 0.00